# Effective Game-Theoretic Motion Planning via Nested Search

### v1

@@ 3 Graph-based setting
To tackle Problem 1, we draw inspiration from previous work on centralized (non-game-theoretic) multi-robot motion, wherein the curse of dimensionality, inherent in exploring the joint state space, is alleviated by first capturing the state space of the individual robots via discrete graphs, and then planning their collective motion on an implicitly-defined tensor graph [solovey2016finding, ShomeSDHB20]. In our setting, however, we utilize specialized constructions of those graphs to accommodate differential constraints (1) and, most importantly, game-theoretic aspects, which have not been considered in this context before. In prepartion for our algorithm, which is described in the next section, we describe here the graph-based setting, as well as a graph-based counterpart of Problem 1.
3.1 Single-robot kinodynamic graphs
@@ 3.2 Multi-robot tensor graph
3.3 Graph-based problem formulation
Next, we adapt Problem 1 to the graph-based setting. Fix a number of steps $n\in\mathbb{N}_{+}$, such that $n\Delta t$ is equal to the planning horizon $T>0$.
In the graph-based setting, a trajectory $\pi^{i}_{n}$ of length $n$, corresponds to a path on the graph $G^{i}$, which consists of a sequence of $n$ edges. That is, given the graph path $e_{1}^{i},\ldots,e_{n}^{i}$, the trajectory $\pi^{i}_{n}$ is the concatenation of the trajectories $\pi^{i}_{e_{1}^{i}},\ldots,\pi^{i}_{e_{n}^{i}}$. We denote by $\Pi^{i}_{n}$ the set of $n$-step trajectories of robot $i$. We similarly define the joint trajectory set $\Pi_{n}:=\times_{i=1}^{n}\Pi^{i}_{n}$. Notice that each joint trajectory $\pi_{n}\in\Pi_{n}$ corresponds to a path on the graph $G$.
We emphasize that the although each trajetory $\pi_{n}\in\Pi_{n}$ is induced by a discrete graph, it represents a continuous and dynamically feasible motion for the robots in $\mathcal{X}$. As such, collision checking and evaluation of trajectory costs with respect to $J,J^{i}$ are performed in the continous domain.
@@ 3.3 Graph-based problem formulation
For a horizon of $n$ steps, a joint discrete trajectory $\pi_{n}\in\Pi_{n}$ is a graph Nash equilibrium (gNE) if, for every robot $i\in[m]$ and every alternative trajectory $\tilde{\pi}^{i}_{n}\in\Pi^{i}_{n}$ (from the same start state), it holds that
$J^{i}(\pi^{i}_{n},\pi^{-i}_{n})\;\leq\;J^{i}(\tilde{\pi}^{i}_{n},\pi^{-i}_{n}).$ (6)
That is, no robot can reduce its cost by unilaterally deviating to another feasible nn–step trajectory.
@@
Fix $n>0$. Given a global cost functional $J:\Pi_{n}\mapsto\mathbb{R}_{\geq 0}$ defined by (4), the objective is to find $\pi^{*}_{n}\in\Pi_{n}$ such that:
  1. 1.
    $\pi^{*}_{n}$ is a graph Nash equilibrium (Def. 2);
  2. 2.
    $\pi^{*}_{n}$ minimizes the global cost: $J(\pi^{*}_{n})$ among all gNEs;
  3. 3.
    $\pi^{*}_{n}$ reaches the joint goal region $\mathcal{X}_{\mathrm{goal}}$ in $n$ steps;
  4. 4.
    $\pi^{*}_{n}$ is collision–free, i.e., all intermediate joint states and the continuous motions along edges respect $\mathcal{X}_{f}$.
Discussion. We discuss the interrelation between Problem 1 and Problem 2. Previous work has shown that for the geometric setting (i.e., without kinodynamic constraints), the graphs $G^{1},\ldots,G^{m}$ can be constructed such that any feasible multi-robot trajectory in the continuous state space of the robot group can be approximated arbitrarily well in the graph $G$ [ShomeSDHB20, Dayan.ea.23]. The extension to the kinodynamic setting can be done via recent analysis techniques for asymptotically optimal sampling-based planning [kleinbort2020refined]. We leave this technical proof for future work. We note that the points highlighted in the discussion in Section 2, are equally applicable to the graph setting.
@@ 4 Algorithm for graph NE
We are ready to describe our algorithmic approach Game-Theoretic Nested Search (GTNS) to tackle Problem 2. The algorithm employs a nested approach, consisting of an outer search (Alg. 1) and an inner search (Alg. 2). Both levels of search are performed by A*-based search algorithms [hart1968formal], albeit on different graphs and for different purposes.
4.1 Outer search
The GTNS algorithm (Alg. 1) explores the implicit tensor-product graph $G$ induced by the per-robot graphs $\{G^{i}\}_{i\in[m]}$ produced by generateRobotRoadmap. These individual graphs can be precomputed once (offline) and reused across scenarios for the same robot platform. By construction, motion along edges of $G$ is dynamically feasible for all robots; along the way, we additionally enforce collision avoidance and the gNE condition.
The search is A-style. It maintains a priority queue OPEN keyed by $J(\pi_{x})+H(x)$ and a predecessor map $P$ for path reconstruction. At each iteration, the algorithm: (i) extracts the minimum-key node $x$ (line 1); (ii) recovers the current best path, denoted $\pi_{x}$, from the root using $P$ (line 1); and (iii) enumerates neighbors $X_{\textup{near}}$ of $x$ in $G$ (line 1). Here $X_{\textup{near}}$ is the tensor product of the per-robot neighbor sets: a joint node $y=(y^{1},\dots,y^{m})\in V$ belongs to $X_{\textup{near}}$ iff for every $i$ there is an edge $(x^{i},y^{i})\in E^{i}$.
For each potential neighbor $x_{\textup{new}}\in X_{\textup{near}}$, the algorithm (a) verifies that the local edge trajectory $\pi_{x\to x_{\textup{new}}}$ is collision-free (line 1); (b) forms the candidate path $\pi_{x_{\textup{new}}}:=\texttt{concatenate}(\pi_{x},\pi_{x\to x_{\textup{new}}})$ (line 1); and (c) checks improvement w.r.t. any incumbent path to $x_{\textup{new}}$ (line 1), where the incumbent is recovered via $P$ (line 1), which is later updated (line 1). It then calls the equilibrium predicate isNashEquilibrium on $\pi_{x_{\textup{new}}}$ (line 1); only if the candidate satisfies the gNE condition is $x_{\textup{new}}$ inserted into OPEN with key $J(\pi_{x_{\textup{new}}})+H(x_{\textup{new}})$. If a goal node is popped and validated, the corresponding trajectory is returned.
@@ 4.1 Outer search
3 $\{G^{i}\}_{i\in[m]}{\leftarrow}\left\{\texttt{generateRobotRoadmap}(f^{i}{,}J^{i}{,}x_{0}^{i}{,}\mathcal{X}_{\text{goal}}^{i})\right\}_{i\in[m]}$
@@ 12       $\pi_{x_{\textup{new}}}\leftarrow\texttt{concatenate}(\pi_{x},\pi_{x\to x_{\textup{new}}})$
13       $\pi_{x_{\textup{new}}}^{\textup{old}}\leftarrow\texttt{trajFromRoot}(P,x_{\textup{new}})$
@@ 17             $P(x_{\textup{new}})\leftarrow x$ $\triangleright$ store parent
4.2 Inner search
isNashEquilibrium is a best-response oracle: it solves a single-robot motion planning problem on $G^{i}$ to see whether any agent has a strictly cheaper unilateral deviation of the same horizon that is dynamically feasible and collision-free w.r.t. others’ fixed paths. For each candidate joint trajectory $\pi_{x_{\textup{goal}}}=(\pi^{i}_{x_{\textup{goal}}^{i}},\pi^{-i}_{x_{\textup{goal}}^{-i}})$ encountered by GTNS, it calls (line 1) isNashEquilibrium (Alg. 2) which runs, for each $i\in[m]$, a standard A search on $G^{i}$ subject to three simple constraints:
  • Terminal-state constraint. The search goal is the specific terminal configuration $x^{i}_{\text{goal}}$ appearing at the end of $\pi^{i}_{x_{\textup{goal}}^{i}}$ (no goal region); see the initialization above Alg. 2.
  • Horizon-length constraint. Only paths of exactly $n$ edges (garnered by the method numEdges) are considered competitive (the same length as $\pi_{x_{\textup{goal}}}$). The code drops nodes once the partial path length $d$ reaches $n$ (the loop continues only while $d<n$).
  • Dynamic-obstacle constraint. Edges for robot $i$ are collision-checked against the fixed, time-parameterized trajectories of the other robots $\pi^{-i}_{x_{\textup{goal}}^{-i}}$; see the collisionFreeEdge method at line 2.
Operationally, for each $i$ the routine initializes OPEN and a predecessor map $P^{i}$, then performs the usual A node selection, neighbor expansion, concatenation, and incumbent comparison on the single-robot graph $G^{i}$ (with key $J^{i}(\cdot)+H^{i}(\cdot)$), exactly mirroring the outer loop but only for agent $i$ and under the constraints above. If it ever reaches $x^{i}_{\text{goal}}$ in exactly $n$ steps with a trajectory $\tilde{\pi}^{i}_{x_{\textup{goal}}^{i}}$ yielding $J^{i}(\tilde{\pi}^{i}_{x_{\textup{goal}}^{i}})<J^{i}(\pi^{i}_{x_{\textup{goal}}^{i}})$, the routine immediately returns False (the candidate joint trajectory does not meet the gNE conditions). If no robot can improve, the routine returns True.
@@ 4.2 Inner search
3 $n\leftarrow\texttt{numEdges}(\pi_{x_{\textup{goal}}})$
@@ 10    $d\leftarrow\texttt{numEdges}(\tilde{\pi}^{i}_{x^{i}})$
19       if $\texttt{collisionFreeEdge}(\tilde{\pi}^{i}_{x^{i}\to x_{\textup{new}}^{i}},\pi^{-i}_{x_{\textup{goal}}^{-i}})$ then
@@ 24             $P^{i}(x_{\textup{new}}^{i})\leftarrow x^{i}$
27             if $x_{\textup{new}}^{i}=x^{i}_{\text{goal}}$ and $d+1=n$ then
28                if $J^{i}(\tilde{\pi}^{i}_{x_{\textup{goal}}^{i}})<J^{i}(\pi^{i}_{x_{\textup{goal}}^{i}})$ then
@@
Algorithm 2 isNashEquilibrium $(\pi_{x_{\textup{goal}}})$
Remark. The pseudocode captures the high-level structure of the GTNS algorithm but omits several implementation optimizations for clarity. The most consequential (inspired by works such as [dellin2016unifying]) is lazy validation of collision and equilibrium checks: every neighbour of $x$ visited in line 1 is pushed to the OPEN list without immediate validation; only when a neighbour is popped from OPEN (i.e., it currently has the lowest cost) do we verify that it is in fact a valid node. This defers the most expensive computations and spends resources only on promising candidates. The implementation also caches edge-collision results (Boolean flags for whether an edge $e\in E$ is in collision) to allow a fast lookup before invoking an expensive check, it precomputes heuristics, and performs early feasibility checks for the query. Additional optimizations include careful memory management and the use of pointers to stable storage.
@@ 5 Theoretical guarantees
Below, we prove that Alg. 1 provides an optimal solution to Problem 2. In preparation, we obtain the following lemma that states that the NE property is monotone, i.e., every subtrajectory of a NE trajectory is also NE.
Throughout this section, given a trajectory $\pi_{n}$ of $n$ edges, the notation $\pi_{n^{\prime}}$ for some $0<n^{\prime}<n$ denotes the subtrajectory of $\pi_{n}$ consisting of the first $n^{\prime}$ edges. Namely, $\pi_{n^{\prime}}$ has a duration of $n^{\prime}\cdot\Delta t$ and $\pi_{n^{\prime}}(\tau)=\pi_{n}(\tau)$ for any $0\leq\tau\leq\Delta t\cdot n^{\prime}$.
@@
If trajectory $\pi_{n}$ is a gNE (Def. 2) then any subtrajectory $\pi_{n^{\prime}}$, where $0\leq n^{\prime}<n$, is a gNE too.
@@
Suppose that $\pi_{n}$ is a gNE, namely for any robot $i\in[m]$, and any alternative trajectory $\tilde{\pi}_{n}^{i}\in\Pi^{i}_{n}$, with the same start and final $\mathcal{X}^{i}$ state, i.e., $\pi^{i}_{n}(0)=\tilde{\pi}^{i}_{n}(0)$, and $\pi^{i}_{n}(n\cdot\Delta t)=\tilde{\pi}^{i}_{n}(n\cdot\Delta t)$, respectively, it holds that
$J^{i}(\pi_{n}^{i},\pi_{n}^{-i})\leq J^{i}(\tilde{\pi}_{n}^{i},\pi_{n}^{-i}).$ (7)
Assume, by contradiction, that the subtrajectory $\pi_{n^{\prime}}$, for some $0<n^{\prime}<n$, is not a gNE. I.e., for some robot $i\in[m]$, there exists $\tilde{\pi}^{i}_{n^{\prime}}\in\Pi^{i}_{n^{\prime}}$ satisfying the inequality
$J^{i}(\pi^{i}_{n^{\prime}},\pi^{-i}_{n^{\prime}})>J^{i}(\tilde{\pi}^{i}_{n^{\prime}},\pi^{-i}_{n^{\prime}}),$
the collision constraints, $\pi^{i}_{n^{\prime}}(0)=\tilde{\pi}^{i}_{n^{\prime}}(0)$, and $\pi^{i}_{n^{\prime}}(n^{\prime}\cdot\Delta t)=\tilde{\pi}^{i}_{n^{\prime}}(n^{\prime}\cdot\Delta t)$.
We now consider the remaining portion of $\pi_{n}$ from the $n^{\prime}+1$ edge, denoted by $\pi_{n^{\prime}+1:n}$. I.e., $\pi_{n^{\prime}+1:n}$ has a duration of $(n-n^{\prime}-1)\cdot\Delta t$ and $\pi_{n^{\prime}+1:n}(\tau)=\pi_{n}(\tau+n^{\prime}\cdot\Delta t)$. Next, denote by $\tilde{\pi}^{i}_{n}$ the trajectory resulting from concatenating $\tilde{\pi}^{i}_{n^{\prime}}$ and $\pi_{n^{\prime}+1:n}$. This new trajectory is valid since $\tilde{\pi}^{i}_{n^{\prime}}$ ends where $\pi_{n^{\prime}+1:n}$ begins, and both those subtrajectories are collision free. Moreover, by exploiting the additivity of the cost function $J^{i}$, we upper bound the $J^{i}$ cost resulting from $\tilde{\pi}^{i}_{n^{\prime}}$:
@@
This contradicts Equation (7). Thus, the gNE property must be monotone. ∎
Combining the above lemma, with the optimality guarantee of A* graph search, which is the backbone of both Alg. 1 and 2, we establish the correctness of GTNS.
@@
Recall that running the A* algorithm returns an optimal solution when running on a search problem $P$ using an admissible heuristic [hart1968formal]. Recall that a search problem is a tuple $P=\langle S,\rm{Succ},s_{\rm begin},S_{\rm goal}\rangle$, where $S$ is a set of (abstract) states, $\rm{Succ}:S\rightarrow 2^{S}$ is a succesor function mapping each state to a set of neighboring states, $s_{\rm begin}\in S$ is the initial state and $S_{\rm goal}\subset S$ a set of goal states.

### v2

@@ 3 Graph-Based setting
To tackle Problem 1, we draw inspiration from previous work on multi-robot motion planning, wherein the curse of dimensionality, inherent in exploring the joint state space, is alleviated by first capturing the state space of the individual robots via discrete graphs, and then planning their collective motion on an implicitly-defined tensor graph [solovey2016finding, ShomeSDHB20, lavalle2002optimal]. In our setting, however, we utilize specialized constructions of those graphs to accommodate differential constraints (1) and, most importantly, game-theoretic aspects. In preparation for our algorithm, which is described in the next section, we describe here the graph-based setting, as well as a graph-based counterpart of Problem 1.
@@
Next, we adapt Problem 1 to the graph-based setting. Fix a number of steps $n\in\mathbb{N}_{+}$, such that $n\Delta t$ is equal to the planning horizon $T>0$. Let a trajectory $\pi^{i}_{n}$ of length $n$ consist of a sequence of $n$ edges on the graph $G_{i}$. That is, given the graph path $e_{1}^{i},\ldots,e_{n}^{i}$, the trajectory $\pi^{i}_{n}$ is the concatenation of the trajectories $\pi^{i}_{e_{1}^{i}},\ldots,\pi^{i}_{e_{n}^{i}}$. We denote by $\Pi^{i}_{n}$ the set of $n$-step trajectories of robot $i$. We similarly define the joint trajectory set $\Pi_{n}:=\times_{i=1}^{n}\Pi^{i}_{n}$. Notice that each joint trajectory $\pi_{n}\in\Pi_{n}$ corresponds to a path on the graph $G$.
@@
For a horizon of $n$ steps, a joint discrete trajectory $\pi_{n}\in\Pi_{n}$ is a graph Nash equilibrium (gNE) if, for every robot $i\in[m]$ and every alternative trajectory $\tilde{\pi}^{i}_{n}\in\Pi^{i}_{n}$ (from the same start state), no robot can reduce its cost by unilaterally deviating to another feasible nn–step trajectory, i.e., $J^{i}(\pi^{i}_{n},\pi^{-i}_{n})\;\leq\;J^{i}(\tilde{\pi}^{i}_{n},\pi^{-i}_{n})$.
@@
Fix $n>0$. Given a global cost $J:\Pi_{n}\mapsto\mathbb{R}_{\geq 0}$, the objective is to find $\pi^{*}_{n}\in\Pi_{n}$ such that: (1) $\pi^{*}_{n}$ is a gNE (Def. 2) that minimizes the global cost: $J(\pi^{*}_{n})$ among all gNEs; (2) $\pi^{*}_{n}$ reaches the joint goal region $\mathcal{X}_{\mathrm{goal}}$ in $n$ steps; (3) $\pi^{*}_{n}$ is collision–free, i.e., all intermediate joint states and the continuous motions along edges respect $\mathcal{X}_{f}$.
Discussion. We discuss the interrelation between Problem 1 and Problem 2. Previous work has shown that for the geometric setting (i.e., without kinodynamic constraints), the graphs $G^{1},\ldots,G^{m}$ can be constructed such that any feasible multi-robot trajectory in the continuous state space of the robot group can be approximated arbitrarily well in the graph $G$ [ShomeSDHB20, Dayan.ea.23]. The extension to the kinodynamic setting can be done via recent analysis techniques for asymptotically optimal sampling-based planning [kleinbort2020refined]. We leave this technical proof for future work. We note that the points highlighted in the discussion in Sec. 2, are equally applicable to the graph setting.
@@ 4 Algorithm for graph NE
We describe our algorithmic approach, Game-Theoretic Nested Search (GTNS), to tackle Problem 2. The algorithm employs a nested approach, consisting of an outer search (Alg. 1) and an inner search (Alg. 2). Both levels of search are performed by A*-based search algorithms [hart1968formal], albeit on different graphs and for different purposes. We describe the high-level structure of the algorithm (see additional implementation optimizations in App. 0.A).
Outer search. The GTNS algorithm (Alg. 1) explores the implicit tensor-product graph $G$ induced by the per-robot graphs $\{G^{i}\}_{i\in[m]}$ each produced by generateRobotRoadmap. These individual graphs can be precomputed once (offline) and reused across scenarios for the same robot platform. By construction, motion along edges of $G$ is dynamically feasible for all robots; we additionally enforce collision avoidance and the gNE condition.
The A* search maintains a priority queue OPEN keyed by $J(\pi_{x})+H(x)$ (where $H(\cdot)$ is a cost-to-go heuristic)) and a predecessor map $P$ for path reconstruction. At each iteration, the algorithm: (i) extracts the minimum-key node $x$ (line 1); (ii) recovers the current best path $\pi_{x}$ from the root using $P$ (line 1); and (iii) enumerates neighbors $X_{\textup{near}}$ of $x$ in $G$ (line 1). Here, a joint node $y=(y^{1},\dots,y^{m})\in V$ belongs to $X_{\textup{near}}$ iff for every $i$ there is an edge $(x^{i},y^{i})\in E^{i}$.
For each potential neighbor $x_{\textup{new}}\in X_{\textup{near}}$, the algorithm (a) verifies that the local edge trajectory $\pi_{x\to x_{\textup{new}}}$ is collision-free (line 1); (b) forms the candidate path $\pi_{x_{\textup{new}}}:=\texttt{concatenate}(\pi_{x},\pi_{x\to x_{\textup{new}}})$ (line 1); and (c) checks improvement w.r.t. any incumbent path to $x_{\textup{new}}$ (line 1), where the incumbent is recovered via $P$ (line 1) and updated (line 1). It then calls the predicate isNashEquilibrium on $\pi_{x_{\textup{new}}}$ (line 1); only if the candidate path satisfies the gNE condition and offers an improved cost over the incumbent path (similar to standard A* dominance checks), is $x_{\textup{new}}$ inserted into OPEN with key $J(\pi_{x_{\textup{new}}})+H(x_{\textup{new}})$. If a goal node is popped and validated, the corresponding trajectory is returned.
@@
3 $\left\{G^{i}\leftarrow\texttt{generateRobotRoadmap}(f^{i}{,}J^{i}{,}x_{0}^{i}{,}\mathcal{X}_{\text{goal}}^{i})\right\}_{i\in[m]}$
@@ 12        $\pi_{x_{\textup{new}}}\leftarrow\texttt{concatenate}(\pi_{x},\pi_{x\to x_{\textup{new}}})$
13        $\pi_{x_{\textup{new}}}^{\textup{old}}\leftarrow\texttt{trajFromRoot}(P,x_{\textup{new}})$ $\triangleright$ incumbent traj. to $x_{\textup{new}}$
@@ 17              $P(x_{\textup{new}})\leftarrow x$ $\triangleright$ store parent
Inner search. isNashEquilibrium is a best-response oracle: it solves a single-robot motion planning problem on $G^{i}$ to see whether any agent has a strictly cheaper unilateral deviation of the same horizon that is dynamically feasible and collision-free w.r.t. others’ fixed paths. For each candidate joint trajectory $\pi_{x_{\textup{cand}}}=(\pi^{i}_{x_{\textup{cand}}^{i}},\pi^{-i}_{x_{\textup{cand}}^{-i}})$ encountered by GTNS, isNashEquilibrium (Alg. 2) runs, for each $i\in[m]$, an A* search on $G^{i}$ subject to three simple constraints: (1) The search goal is set (line 2) to the specific terminal configuration $x_{\textup{cand}}^{i}$ appearing at the end of $\pi^{i}_{x_{\textup{cand}}^{i}}$ (no goal region); (2) Only paths of exactly $n$ edges (garnered by the method numEdges) are considered competitive (the same length as $\pi_{x_{\textup{cand}}}$). Nodes are only considered until the partial path length $d$ reaches $n$; (3) Edges for robot $i$ are collision-checked (collisionFreeEdge, line 2) against the fixed, time-parameterized trajectories of the other robots $\pi^{-i}_{x_{\textup{cand}}^{-i}}$.
Operationally, for each $i$ the routine initializes OPEN and a predecessor map $P^{i}$, then performs the usual A node selection, neighbor expansion, concatenation, and incumbent comparison on the single-robot graph $G^{i}$ (with key $J^{i}(\cdot)+H^{i}(\cdot)$), exactly mirroring the outer loop but only for agent $i$ and under the constraints above. If it ever reaches $x_{\textup{cand}}^{i}$ in exactly $n$ steps with a trajectory $\tilde{\pi}^{i}_{x_{\textup{cand}}^{i}}$ yielding $J^{i}(\tilde{\pi}^{i}_{x_{\textup{cand}}^{i}})<J^{i}(\pi^{i}_{x_{\textup{cand}}^{i}})$, the routine immediately returns False (the candidate joint trajectory does not meet the gNE conditions). If no robot can improve, the routine returns True.
@@
3 $n\leftarrow\texttt{numEdges}(\pi_{x_{\textup{cand}}})$
@@ 10     $d\leftarrow\texttt{numEdges}(\tilde{\pi}^{i}_{x^{i}})$
19        if $\texttt{collisionFreeEdge}(\tilde{\pi}^{i}_{x^{i}\to x_{\textup{new}}^{i}},\pi^{-i}_{x_{\textup{cand}}^{-i}})$ then
@@ 24              $P^{i}(x_{\textup{new}}^{i})\leftarrow x^{i}$
27             if $x_{\textup{new}}^{i}=x_{\textup{cand}}^{i}$ and $d+1=n$ then
28                 if $J^{i}(\tilde{\pi}^{i}_{x_{\textup{cand}}^{i}})<J^{i}(\pi^{i}_{x_{\textup{cand}}^{i}})$ then
@@
Algorithm 2 isNashEquilibrium $(\pi_{x_{\textup{cand}}})$
@@ 5 Theoretical guarantees
We prove that Alg. 1 provides an optimal solution to Problem 2. In preparation, we obtain the following lemma that states that the NE property is monotone, i.e., every subtrajectory of a NE trajectory is also NE. Given a trajectory $\pi_{n}$ of $n$ edges, the notation $\pi_{n^{\prime}}$ for some $0<n^{\prime}<n$ denotes the subtrajectory of $\pi_{n}$ consisting of the first $n^{\prime}$ edges. Namely, $\pi_{n^{\prime}}$ has a duration of $n^{\prime}\cdot\Delta t$ and $\pi_{n^{\prime}}(\tau)=\pi_{n}(\tau)$ for any $0\leq\tau\leq\Delta t\cdot n^{\prime}$.
@@
If a trajectory $\pi_{n}$ is a gNE (Def. 2) then any subtrajectory $\pi_{n^{\prime}}$, where $0\leq n^{\prime}<n$, is a gNE too.
@@
First, recall that we consider additive cost functions $J,\{J^{i}\}$. Suppose that $\pi_{n}$ is a gNE, namely for any robot $i\in[m]$, and any alternative trajectory $\tilde{\pi}_{n}^{i}\in\Pi^{i}_{n}$, with the same start and final $\mathcal{X}^{i}$ state, i.e., $\pi^{i}_{n}(0)=\tilde{\pi}^{i}_{n}(0)$, and $\pi^{i}_{n}(n\cdot\Delta t)=\tilde{\pi}^{i}_{n}(n\cdot\Delta t)$, respectively, it holds that $J^{i}(\pi_{n}^{i},\pi_{n}^{-i})\leq J^{i}(\tilde{\pi}_{n}^{i},\pi_{n}^{-i})$.
Assume, by contradiction, that the subtrajectory $\pi_{n^{\prime}}$, for some $0<n^{\prime}<n$, is not a gNE: for some robot $i\in[m]$, there exists $\tilde{\pi}^{i}_{n^{\prime}}\in\Pi^{i}_{n^{\prime}}$ satisfying the collision constraints, $\pi^{i}_{n^{\prime}}(0)=\tilde{\pi}^{i}_{n^{\prime}}(0)$, and $\pi^{i}_{n^{\prime}}(n^{\prime}\cdot\Delta t)=\tilde{\pi}^{i}_{n^{\prime}}(n^{\prime}\cdot\Delta t)$ and inequality $J^{i}(\pi^{i}_{n^{\prime}},\pi^{-i}_{n^{\prime}})>J^{i}(\tilde{\pi}^{i}_{n^{\prime}},\pi^{-i}_{n^{\prime}})$. Consider the remaining portion of $\pi_{n}$ from the $n^{\prime}+1$ edge, denoted by $\pi_{n^{\prime}+1:n}$. I.e., $\pi_{n^{\prime}+1:n}$ has a duration of $(n-n^{\prime}-1)\cdot\Delta t$ and $\pi_{n^{\prime}+1:n}(\tau)=\pi_{n}(\tau+n^{\prime}\cdot\Delta t)$. Next, denote by $\tilde{\pi}^{i}_{n}$ the concatenation of $\tilde{\pi}^{i}_{n^{\prime}}$ and $\pi_{n^{\prime}+1:n}$. This new trajectory is valid as $\tilde{\pi}^{i}_{n^{\prime}}$ ends where $\pi_{n^{\prime}+1:n}$ begins, and both those subtrajectories are collision free. By exploiting the additivity of the cost function $J^{i}$, we upper bound the $J^{i}$ cost resulting from $\tilde{\pi}^{i}_{n^{\prime}}$:
@@
This contradicts $\tilde{\pi}_{n}$ being a gNE. Thus, the gNE property is monotone. ∎
Combining the above lemma, with the optimality guarantee of A*, which is the backbone of both Alg. 1 and 2, we establish the correctness of GTNS.
@@
The A* algorithm returns an optimal solution when running on a search problem $P$ using an admissible heuristic [hart1968formal]. Recall that a search problem is a tuple $P=\langle S,\rm{Succ},s_{\rm begin},S_{\rm goal}\rangle$, where $S$ is a set of (abstract) states, $\rm{Succ}:S\rightarrow 2^{S}$ is a succesor function mapping each state to a set of neighboring states, $s_{\rm begin}\in S$ is the initial state and $S_{\rm goal}\subset S$ a set of goal states.
@@
Time complexity. The branching factor of node expansion in GTNS grows exponentially with the number of robots. Even so, our approach efficiently copes with realistic problem settings in a few seconds (see Sec. 6), due to the aggressive pruning of isNashEquilibrium, which substantially reduces the search space, as compared to the non-game-theoretic and centralized multi-robot motion planning setting [WagnerC15]. We discuss additional speed-up techniques in Sec. 7.